OpenStudy (anonymous):

select the ordered pair that is the solution of this equation -7x-4y=6

OpenStudy (zpupster):

are there choices??

OpenStudy (zpupster):

let x be 0 then -4y = 6 y = 6/-4 or -3/2 let y be 0 and -7x = 6 x=-6/7 (x,y) or (-6/7, -3/2)

OpenStudy (zpupster):

oops the ordered pairs would be (0,-3/2) and( -6/7, 0)

OpenStudy (anonymous):

your right the ordered pairs would be (0,-3/2) and( -6/7, 0)

OpenStudy (anonymous):

the general solution for x=6+4n , y=-12-7n

OpenStudy (swag):

Hi, I can see that you are a new member here at OpenStudy so, I would first like to say Welcome to OpenStudy, I would like to point you to the chat pods these chat pods are where you can make new friends and talk to new people just like you, I would also like to emphasize our "NON-CHEATING POLICY" we ask here at OpenStudy that you do NOT post exam/Test questions if you are caught doing this we will notify your current school I would also like to ask you to check out the OpenStudy "Code of Conduct" http://openstudy.com/code-of-conduct Please read this carefully and thoroughly. If you need any help feel free to message me or any other ambassador. Welcome to the OpenStudy Community! Ambassador SWAG